OpenStudy (anonymous):

[1 -2 3 x] For what value of x in this matrix is there no inverse?

OpenStudy (mathmale):

Hint: consider the matrix\[\left[ \left[\begin{matrix}a & b \\ c & d\end{matrix}\right] \right]\]

OpenStudy (mathmale):

If ab = cd, the matrix does not have an inverse (due to division by zero).

OpenStudy (mathmale):

Use this fact to determine the value of x.

OpenStudy (anonymous):

I'm sorry, I still don't understand.

OpenStudy (anonymous):

u still need help?

OpenStudy (anonymous):

yes please

OpenStudy (anonymous):

ok remember a metrix does not have an inverse if the determinant is equal to 0 right?

OpenStudy (anonymous):

Right

OpenStudy (anonymous):

now to elaborate more on what mathmale had stated above.. If ab = cd, the matrix does not have an inverse. in your case 1 times x=? 3 times -2=-6 so then 1 times -6=-6

OpenStudy (anonymous):

now try to find the determinant what do you get

OpenStudy (anonymous):

so x=-6 ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

u know how to find the determinant?

OpenStudy (anonymous):

yes I do, I found the determinant is 0 so that means x=-6. So the main goal is to get the determinant to equal 0?

OpenStudy (anonymous):

yes

OpenStudy (mathmale):

Note that "ad=bc" is equivalent to "the determinant is 0."

OpenStudy (anonymous):

okay thank you!

OpenStudy (mathmale):

magepker728, your clarifications were very good.

OpenStudy (anonymous):

thanks @mathmale